# STARFormer: A Novel Spatio-Temporal Aggregation Reorganization Transformer of FMRI for Brain Disorder Diagnosis

### v1

@@ 3.3.1 Window Tokens and Extended Window Tokens
In the temporal branch, the arranged fMRI time series data $S^{\prime}$ are segmented into equally sized window tokens $\{x_{1}^{l},x_{2}^{l},...,x_{g}^{l}\}$ along the temporal dimension by the variable window, where $g$ is the number of window tokens and $l$ is the layer of the temporal branch. The length of each window token $x$ is $w=m/g$. The extended window token $y_{i}$ is formed by extending $x_{i}$ at both the start and the end of the temporal dimension by a length of $w/2$, respectively. The length of $y_{i}$ is $2w$. To ensure that the lengths of the extended window tokens are consistent, padding of length $w/2$ is added to the start of the first window token $x_{1}$ and the end of the last window token $x_{g}$ to form the extended window tokens $y_{1}$ and $y_{g}$, while the overall length of the time series extends to $m+w$. It is intended that the positions of this padding do not participate in the subsequent backpropagation process, so masks will be used to cover the padded positions during training to avoid affecting the performance of the model.
@@ 3.3.2 Transformer Block
Performing global self-attention requires calculating the relationships between each token and all other tokens, which undoubtedly involves high computational complexity. However, using the extended window in the attention computation allows the representation of local information to no longer be limited to its own window tokens. This facilitates the interaction of information between window tokens and reduces computational complexity. To better capture specific positional information between window tokens, bias is incorporated into the attention computation to adjust the attention weight matrix [37, 38]:
@@
where $\mathbf{B}$ is a learnable positional bias matrix and $d$ is the feature dimension of the attention heads. $\mathbf{B}$ represents the positions of the window tokens relative to all the tokens in the receptive field, including both the window tokens and the extended window tokens. The FF layer applies a nonlinear transformation to the input and utilizes Dropout to prevent model overfitting. The activation function used is GELU (Gaussian error linear unit) [39].
@@ 3.3.3 Variable Window
for $l=1,2,...,6$ do # $l:$ l-th layer of temperal branch
@@ 4.1 Datasets
In this study, we conducted experiments using two public fMRI datasets: ABIDE-I and ADHD-200. The ABIDE-I dataset was compiled by collaboration between 17 international imaging sites, openly sharing 871 valid fMRI samples from 403 individuals with ASD and 468 typically developing controls [40]. The ADHD-200 dataset was collected from 8 international imaging sites, openly sharing 947 valid fMRI samples from 362 children and adolescents with ADHD and 585 typically developing controls [41].
The preprocessed fMRI datasets are available in C-PAC of ABIDE-I and Athena of ADHD-200 [42, 43]. Specifically, the preprocessing steps include voxel intensity normalization, motion correction, and slice timing correction. The fMRI images are then co-registered to their corresponding anatomical images and normalized to MNI152 space. Finally, the mean time series is extracted from each ROI for each subject on the basis of the specified atlas.
@@ 4.2 Experimental Process and Details
We randomly selected 10% of patient samples from the dataset as input into the ROI spatial structure analysis module. For each subject in the dataset, the fMRI time series was randomly cropped to 128 samples along the temporal dimension to maximize the retention of the sample information. ROI parcellation was determined using two public brain atlases: the Schaefer atlas [44] and the AAL atlas [45]. For the Schaefer atlas, we selected the scale of 400 ROIs across seven intrinsic connectivity networks. The AAL atlas partitions the brain anatomically into 116 ROIs. The STARFormer model was trained for 100 epochs with a batch size of 128 and a dropout rate of 0.5. Eight attention heads were specified, each with 16 dimensions. For training on ABIDE-I dataset, the initial learning rate was set to 5e-5, with a maximum of 1e-4, and gradually reduced to 1e-5. For training on ADHD-200 dataset, the initial learning rate was set to 1e-5, with a maximum of 5e-5, and finally reduced to 1e-6.
The experiments were conducted by using PyTorch based on an NVIDIA RTX 2080Ti GPU. We evaluated the performance of the proposed model using a 10-fold cross-validation, and divided the data into a non-overlapping training set (80%), a validation set (10%), and a test set (10%). The model was trained using the Adam optimizer, with cross-entropy loss as the loss function. To comprehensively evaluate the performance of the model, we employed four commonly used metrics, including accuracy (Acc), precision (Prec), recall (Rec), and area under curve (AUC).
4.3 Competing Methods
Advanced traditional baselines (SVM[46], LSTM[27], BrainNetCNN[47]), transformer-based models (SwinTransformer[37], BolT[28], Com-BrainTF[48]), and graph neural networks (GNN) (MAHGCN[49], RGTNet[50], PLSNet[51]) were selected for comparison with STARFormer. The architectures, loss functions, and learning rate schedulers for each competing method were adopted from their original papers and subsequently fine-tuned in the experimentation to ensure optimal and competitive performance.
@@ 4.3 Competing Methods
Schaefer Model Acc Prec Rec AUC Acc Prec Rec AUC
SVM[46] 65.72±4.11 53.97±7.33 66.26±6.68 71.33±4.62 59.76±2.68 59.34±5.32 59.61±5.43 60.62±4.66
LSTM[27] 66.78±4.25 64.05±9.93 65.81±7.62 70.25±3.85 65.06±3.96 62.92±4.30 60.48±6.38 65.56±5.57
Traditional BrainNetCNN[47] 67.34±4.22 63.82±8.73 65.50±4.63 73.15±5.62 66.78±3.90 62.62±4.66 62.82±8.33 62.22±4.13
SwinT[37] 69.79±4.59 60.75±7.60 68.27±6.56 74.13±4.15 68.56±4.74 65.25±4.36 66.40±9.70 73.85±4.59
BolT[28] 71.28±4.62 69.85±4.94 71.32±4.35 77.46±3.44 70.82±3.57 68.04±4.18 71.51±3.27 73.36±3.86
Transformer Com-BrainTF[48] 72.75±4.56 70.65±4.54 78.43±4.63 77.93±3.02 68.94±2.68 67.36±4.23 72.92±3.88 73.54±3.68
MAHGCN[49] 73.12±3.63 71.05±5.38 72.02±4.14 72.07±3.03 70.76±4.63 69.95±3.38 71.08±2.14 74.25±3.41
RGTNet[50] 74.43±4.82 73.67±4.38 75.28±4.06 74.55±2.73 72.03±3.11 70.14±3.81 71.33±4.45 74.28±2.82
GNN PLSNet[51] 75.17±4.62 75.91±3.65 79.82±5.83 77.03±2.17 72.53±4.27 73.69±3.82 72.42±3.22 78.02±2.87
Ours STARFormer 77.57±3.70 76.98±3.27 84.38±3.50 78.29±3.23 74.12±2.47 73.37±3.07 73.54±3.27 78.81±2.18
@@
AAL Model Acc Prec Rec AUC Acc Prec Rec AUC
SVM[46] 63.74±3.99 52.35±7.18 63.60±6.35 69.90±4.53 57.96±2.60 56.96±5.16 57.82±5.16 59.40±4.52
LSTM[27] 64.77±4.12 62.12±9.73 63.17±7.24 68.84±3.77 62.10±3.84 60.69±4.17 58.66±6.06 64.24±5.40
Traditional BrainNetCNN[47] 65.31±4.09 60.90±8.56 62.88±4.40 70.68±5.51 61.77±3.78 60.11±4.52 60.63±7.91 60.77±4.01
SwinT[37] 65.69±4.45 58.92±7.45 60.53±0.53 72.64±4.07 66.50±4.60 62.64±4.23 65.22±9.22 66.29±4.45
BolT[28] 69.41±2.15 68.52±4.07 66.49±4.22 73.30±3.51 67.66±3.46 68.15±4.05 68.36±5.01 69.83±3.74
Transformer Com-BrainTF[48] 70.56±4.42 68.53±4.41 77.21±4.25 75.37±2.96 66.87±2.60 69.74±4.10 70.73±3.59 72.06±3.51
MAHGCN[49] 71.31±3.52 69.40±3.27 70.09±3.93 71.11±2.97 69.69±2.49 71.36±3.28 70.19±4.23 71.86±3.31
RGTNet[50] 73.19±3.68 71.45±3.29 73.22±3.86 72.32±2.68 70.06±2.96 70.02±3.65 73.41±3.70 74.75±2.74
GNN PLSNet[51] 72.97±3.48 70.87±3.54 76.62±3.54 75.46±3.25 70.53±3.14 71.81±2.03 72.97±4.86 75.45±2.78
Ours STARFormer 75.19±2.93 73.96±2.13 79.27±3.42 75.91±2.85 72.92±2.40 72.59±2.02 73.12±3.23 76.39±2.45
@@ 5.1 Comparative Studies
  Variable Window ABIDE-I ADHD-200
@@ 6.1 Interpretability Analysis
As shown in Fig. 4, we present the top 5% most influential ROIs for diagnosing ASD and ADHD. A manual review confirmed that all identified ROIs are consistent with the previous literature, linking them with the neural manifestations of ASD and ADHD. For instance, functional abnormalities in the parahippocampal gyrus may impair the ability of individuals with ASD to adapt or recall relevant information in social settings, affecting their social behavior [18]. Siminarly, functional abnormalities in the amygdala can lead to anxiety or avoidance in social situations for those with ASD [52]. In addition, functional abnormalities in the precentral gyrus have been associated with hyperactivity symptoms in individuals with ADHD [53], while abnormalities in the inferior temporal gyrus can result in attention deficits or distractibility during complex visual tasks [54]. Overall, this demonstrates that the STARFormer effectively captures brain activation patterns in both healthy individuals and those with brain disorders.
@@ 6.2 Limitation and Future Work
While our proposed STARFormer shows significant improvement over existing computer-aided diagnosis methods for brain disorders, several issues should be considered in future work. First, we used node centrality measures to extract spatial information for ROIs rather than directly using the topological information of the brain network. Future research could explore using graph encoding to model brain topology. Second, auxiliary information about patients, such as personal details or scanning protocols, was not included as input to the model. Recent studies suggest that these phenotypic data complement imaging data and may improve the diagnosis of brain disorders[55]. Therefore, it is reasonable to expect that integrating phenotypic information could further improve the classification performance of STARFormer. Lastly, considering the challenges in data acquisition in clinical settings, many subjects will have partially labeled fMRI data. Thus, strategies using semi-supervised or unsupervised learning could be considered for training.

### v2

@@ 3.3.1 Window Tokens and Extended Window Tokens
In the temporal branch, the arranged fMRI time series data $S^{\prime}$ are segmented into equally sized window tokens $\{x_{1}^{l},x_{2}^{l},...,x_{g}^{l}\}$ along the temporal dimension by the variable window, where $g$ is the number of window tokens and $l$ is the layer of the temporal branch. The length of each window token $x$ is $w=m/g$. The extended window token $y_{i}$ is formed by extending $x_{i}$ at both the start and the end of the temporal dimension by a length of $w/2$, respectively. The length of $y_{i}$ is $2w$. To ensure that the lengths of the extended window tokens are consistent, padding of length $w/2$ is added to the start of the first window token $x_{1}$ and the end of the last window token $x_{g}$ to form the extended window tokens $y_{1}$ and $y_{g}$, while the overall length of the time series extends to $m+w$. It is intended that the positions of this padding do not participate in the subsequent backpropagation process, so zero vectors of length w/2 will be used to cover the padded positions during training to avoid affecting the performance of the model.
@@ 3.3.2 Transformer Block
Performing global self-attention requires calculating the relationships between each token and all other tokens, which undoubtedly involves high computational complexity. However, using the extended window in the attention computation allows the representation of local information to no longer be limited to its own window tokens. This facilitates the interaction of information between window tokens and reduces computational complexity. To better capture specific positional information between window tokens, bias is incorporated into the attention computation to adjust the attention weight matrix (liu2021swin, ; yang2021focal, ):
@@
where $\mathbf{B}$ is a learnable positional bias matrix and $d$ is the feature dimension of the attention heads. $\mathbf{B}$ represents the positions of the window tokens relative to all the tokens in the receptive field, including both the window tokens and the extended window tokens. The FF layer applies a nonlinear transformation to the input and utilizes Dropout to prevent model overfitting. The activation function used is GELU (Gaussian error linear unit) (hendrycks2016gaussian, ).
@@ 3.3.3 Variable Window
To quantify the computational benefits of our cross-window attention mechanism, we provide a rigorous complexity comparison with global self-attention. For a sequence of length $m$ with hidden dimension $n$, global self-attention requires computing attention weights between all token pairs, yielding time complexity $\mathcal{O}(m^{2}n)$. In contrast, our cross-window attention divides the sequence into $g$ windows of size $w=m/g$, where each window processes extended tokens of size $2w$ to capture cross-window dependencies. This results in time complexity $\mathcal{O}(g\times(2w)^{2}n)=\mathcal{O}(4m^{2}n/g)$, achieving a computational reduction factor of $g/4$ compared to global self-attention.
@@
for l=1,2,,6l=1,2,...,6 do # $l:$ l-th layer of temperal branch
@@ 4.1 Datasets
In this study, we conducted experiments using two public fMRI datasets: ABIDE-I and ADHD-200. The ABIDE-I dataset was compiled by collaboration between 17 international imaging sites, openly sharing 871 valid fMRI samples from 403 individuals with ASD and 468 typically developing controls (craddock2013neuro, ). The ADHD-200 dataset was collected from 8 international imaging sites, openly sharing 947 valid fMRI samples from 362 children and adolescents with ADHD and 585 typically developing controls (bellec2017neuro, ).
The preprocessed fMRI datasets are available in C-PAC of ABIDE-I and Athena of ADHD-200 (cpac, ; bellec2017neuro, ). Specifically, the preprocessing steps include voxel intensity normalization, motion correction, and slice timing correction. The fMRI images are then co-registered to their corresponding anatomical images and normalized to MNI152 space. Finally, the mean time series is extracted from each ROI for each subject on the basis of the specified atlas.
@@ 4.2 Experimental Process and Details
We randomly selected 10% of patient samples from the dataset as input into the ROI spatial structure analysis module. For each subject in the dataset, the fMRI time series was randomly cropped to 128 samples along the temporal dimension to maximize the retention of the sample information. We chose the temporal length by considering the average duration of our samples to preserve as much information as possible. In addition, since each sample must be divided into an integer number of window tokens, we ultimately set the temporal dimension to 128. ROI parcellation was determined using two public brain atlases: the Schaefer atlas (schaefer2018local, ) and the AAL atlas (tzourio2002automated, ). The Schaefer atlas parcels the brain based on functional connectivity, emphasizing functional coherence, while the AAL atlas is organized by anatomical structure, facilitating comparison with traditional neuroanatomical studies. Using both atlases allows us to test our model’s robustness under different spatial parcellation schemes. These two atlases are widely used in neuroimaging, making it easier to compare our results with existing work and to enhance interpretability and generalizability. For the Schaefer atlas, we selected the scale of 400 ROIs across seven intrinsic connectivity networks. The AAL atlas partitions the brain anatomically into 116 ROIs. The STARFormer model was trained for 100 epochs with a batch size of 128 and a dropout rate of 0.5. Eight attention heads were specified, each with 16 dimensions. For training on ABIDE-I dataset, the initial learning rate was set to 5e-5, with a maximum of 1e-4, and gradually reduced to 1e-5. For training on ADHD-200 dataset, the initial learning rate was set to 1e-5, with a maximum of 5e-5, and finally reduced to 1e-6.
@@
Model Acc Prec Rec AUC Acc Prec Rec AUC
SVM (abraham2017deriving, ) 65.72±4.11 53.97±7.33 66.26±6.68 71.33±4.62 59.76±2.68 59.34±5.32 59.61±5.43 60.62±4.66
LSTM (dvornek2017lstm, ) 66.78±4.25 64.05±9.93 65.81±7.62 70.25±3.85 65.06±3.96 62.92±4.30 60.48±6.38 65.56±5.57
BrainNetCNN (kawahara2017brainnetcnn, ) 67.34±4.22 63.82±8.73 65.50±4.63 73.15±5.62 66.78±3.90 62.62±4.66 62.82±8.33 62.22±4.13
SwinT (liu2021swin, ) 69.79±4.59 60.75±7.60 68.27±6.56 74.13±4.15 68.56±4.74 65.25±4.36 66.40±9.70 73.85±4.59
BolT (bedel2023bolt, ) 71.28±4.62 69.85±4.94 71.32±4.35 77.46±3.44 70.82±3.57 68.04±4.18 71.51±3.27 73.36±3.86
Com-BrainTF (bannadabhavi2023community, ) 72.75±4.56 70.65±4.54 78.43±4.63 77.93±3.02 68.94±2.68 67.36±4.23 72.92±3.88 73.54±3.68
MAHGCN (liu2023mahgcn, ) 73.12±3.63 71.05±5.38 72.02±4.14 72.07±3.03 70.76±4.63 69.95±3.38 71.08±2.14 74.25±3.41
RGTNet (wang2024rgtnet, ) 74.43±4.82 73.67±4.38 75.28±4.06 74.55±2.73 72.03±3.11 70.14±3.81 71.33±4.45 74.28±2.82
PLSNet (wang2023plsnet, ) 75.17±4.62 75.91±3.65 79.82±5.83 77.03±2.17 72.53±4.27 73.69±3.82 72.42±3.22 78.02±2.87
STARFormer 77.57±3.70 76.98±3.27 84.38±3.50 78.29±3.23 74.12±2.47 73.37±3.07 73.54±3.27 78.81±2.18
@@
Model Acc Prec Rec AUC Acc Prec Rec AUC
SVM (abraham2017deriving, ) 63.74±3.99 52.35±7.18 63.60±6.35 69.90±4.53 57.96±2.60 56.96±5.16 57.82±5.16 59.40±4.52
LSTM (dvornek2017lstm, ) 64.77±4.12 62.12±9.73 63.17±7.24 68.84±3.77 62.10±3.84 60.69±4.17 58.66±6.06 64.24±5.40
BrainNetCNN (kawahara2017brainnetcnn, ) 65.31±4.09 60.90±8.56 62.88±4.40 70.68±5.51 61.77±3.78 60.11±4.52 60.63±7.91 60.77±4.01
SwinT (liu2021swin, ) 65.69±4.45 58.92±7.45 60.53±0.53 72.64±4.07 66.50±4.60 62.64±4.23 65.22±9.22 66.29±4.45
BolT (bedel2023bolt, ) 69.41±2.15 68.52±4.07 66.49±4.22 73.30±3.51 67.66±3.46 68.15±4.05 68.36±5.01 69.83±3.74
Com-BrainTF (bannadabhavi2023community, ) 70.56±4.42 68.53±4.41 77.21±4.25 75.37±2.96 66.87±2.60 69.74±4.10 70.73±3.59 72.06±3.51
MAHGCN (liu2023mahgcn, ) 71.31±3.52 69.40±3.27 70.09±3.93 71.11±2.97 69.69±2.49 71.36±3.28 70.19±4.23 71.86±3.31
RGTNet (wang2024rgtnet, ) 73.19±3.68 71.45±3.29 73.22±3.86 72.32±2.68 70.06±2.96 70.02±3.65 73.41±3.70 74.75±2.74
PLSNet (wang2023plsnet, ) 72.97±3.48 70.87±3.54 76.62±3.54 75.46±3.25 70.53±3.14 71.81±2.03 72.97±4.86 75.45±2.78
STARFormer 75.19±2.93 73.96±2.13 79.27±3.42 75.91±2.85 72.92±2.40 72.59±2.02 73.12±3.23 76.39±2.45
The experiments were conducted by using PyTorch based on an NVIDIA RTX 2080Ti GPU. We evaluated the performance of the proposed model using a 10-fold cross-validation, and divided the data into a non-overlapping training set (80%), a validation set (10%), and a test set (10%). All random sampling in our study was performed using simple random sampling without any stratification. The model was trained using the Adam optimizer, with cross-entropy loss as the loss function. To comprehensively evaluate the performance of the model, we employed four commonly used metrics, including accuracy (Acc), precision (Prec), recall (Rec), and area under curve (AUC).
4.3 Competing Methods
This study selects three types of baseline models for comparison with STARFormer, namely advanced traditional baselines, transformer-based models, and graph neural networks. Transformer-based models excel at analyzing time series data, while graph neural networks focus on capturing spatial information and modeling spatial structures. All these types of models have been applied in fMRI research, demonstrating strong performance and good generalizability.The architectures, loss functions, and learning rate schedulers for each competing method were adopted from their original papers and subsequently fine-tuned in the experimentation to ensure optimal and competitive performance.
Figure 3: ROC curves for comparing the performance metrics of different methods on ABIDE-I and ADHD-200 datasets using Schaefer and AAL atlases.
@@ 5.1 Comparative Studies
The ROC curve in Fig.3 shows the trade-off relationship between the true positive rate and the false positive rate of different methods. The AUC value further quantifies the overall discriminative ability of the classifier. The results show that STARFormer exhibits excellent ROC performance under all dataset configurations, indicating that this method has outstanding classification discrimination ability and clinical application potential.
@@
Figure 5: Confusion matrices for all classification methods on ABIDE-1 (a) and ADHD-200 (b) datasets using Schaefer atlases.
Figure 6: Classification accuracy with 95% confidence intervals for different methods on ABIDE-1 (a) and ADHD-200 (b) datasets using Schaefer atlases.
Figure 7: Cohen’s d effect sizes comparing different classification methods with STARFormer on ABIDE-1 (a) and ADHD-200 (b) datasets using Schaefer atlases.
Fig.5 summarizes the classification outcomes for all methods on both datasets via confusion matrices. The stability and reliability of STARFormer are further demonstrated in Fig.6, where the 95% confidence intervals for accuracy on both ABIDE-1 and ADHD-200 datasets are notably narrow, indicating consistent performance across different cohorts. Regarding effect size analysis, Fig.7 reveals that STARFormer achieves substantial improvements over traditional baseline methods, with effect size differences exceeding 0.2, which indicates significant classification advantages. Meanwhile, the performance gap between Transformer-based models and graph neural network methods shows more moderate differences, with effect sizes typically around 0.1, suggesting that both advanced approaches demonstrate comparable capabilities while STARFormer maintains a competitive edge.
@@
  Window ABIDE-I ADHD-200
@@ 5.2 Ablation Studies
We implement t-SNE-based feature visualization with silhouette coefficient quantification for each module. Fig.8 demonstrates how the ROI Spatial Structure Analysis Module enhances spatial feature separability through EC reorganization. Fig.9 shows how the Temporal Feature Reorganization Module improves temporal pattern discrimination via variable window mechanisms. Fig.10 illustrates how the Spatio-Temporal Feature Fusion Module creates superior integrated representations by combining dual-branch features.
Figure 8: ROI Spatial Structure Analysis Module enhances spatial feature separability through EC reorganization.
Figure 9: Temporal Feature Reorganization Module. (a) Variable window attention vs. global attention comparison. (b) Temporal feature enhancement via t-SNE visualization.
Figure 10: Spatio-Temporal Feature Fusion Module. (a) Dual-branch integration process from spatial and temporal features to fused representation. (b) Feature fusion effectiveness.
@@ 5.3 Window Setting
Figure 12: Temporal attention analysis showing ROI importance across temporal branch.
@@ 6.1 Interpretability Analysis
We implement a systematic ROI identification procedure using multi-head attention weight analysis. Fig.12 shows the temporal attention analysis, where we extract attention weights from STARFormer’s temporal branch architecture across different temporal windows and brain regions organized by functional networks. This heatmap visualization reveals which ROIs receive the highest attention weights during classification, providing a systematic ranking of ROI importance.
As shown in Fig.13, we present the top 5% most influential ROIs for diagnosing ASD and ADHD. A manual review confirmed that all identified ROIs are consistent with the previous literature, linking them with the neural manifestations of ASD and ADHD. For instance, functional abnormalities in the parahippocampal gyrus may impair the ability of individuals with ASD to adapt or recall relevant information in social settings, affecting their social behavior (monk2009abnormalities, ). Siminarly, functional abnormalities in the amygdala can lead to anxiety or avoidance in social situations for those with ASD (kleinhans2009reduced, ). In addition, functional abnormalities in the precentral gyrus have been associated with hyperactivity symptoms in individuals with ADHD (lei2015functional, ), while abnormalities in the inferior temporal gyrus can result in attention deficits or distractibility during complex visual tasks (kobel2010structural, ). Overall, this demonstrates that the STARFormer effectively captures brain activation patterns in both healthy individuals and those with brain disorders.
@@ 6.2 Limitation and Future Work
While our proposed STARFormer shows significant improvement over existing computer-aided diagnosis methods for brain disorders, several issues should be considered in future work. First, we used node centrality measures to extract spatial information for ROIs rather than directly using the topological information of the brain network. Future research could explore using graph encoding to model brain topology. Second, auxiliary information about patients, such as personal details or scanning protocols, was not included as input to the model. Recent studies suggest that these phenotypic data complement imaging data and may improve the diagnosis of brain disorders (Zhang2024Preserving, ). Therefore, it is reasonable to expect that integrating phenotypic information could further improve the classification performance of STARFormer. Lastly, considering the challenges in data acquisition in clinical settings, many subjects will have partially labeled fMRI data. Thus, strategies using semi-supervised or unsupervised learning could be considered for training.